OpenStudy (anonymous):

I REALLY NEED HELP!! what is the simplified form of the expression sqrt 16c^4- sqrt c^2+3 sqrt c^2+ sqrt 9c^3

OpenStudy (mathstudent55):

Is this the problem? \( \sqrt {16c^4} - \sqrt{ c^2}+3 \sqrt {c^2}+ \sqrt {9c^3}\)

OpenStudy (anonymous):

no the last sqrt is sqrt 9c^2

OpenStudy (anonymous):

thank you

OpenStudy (mathstudent55):

\(\sqrt {16c^4} - \sqrt{ c^2}+3 \sqrt {c^2}+ \sqrt {9c^2}\) In general: \( \sqrt{a^2} = |a| \) \( \sqrt{a^4} = a^2 \)

OpenStudy (anonymous):

yes that is right! so do u think you can help me with it?

OpenStudy (anonymous):

im taking a module test on FLVS and I seem to be having a hard time with it

OpenStudy (mathstudent55):

Have you been told that c is positive or nonnegative?

OpenStudy (anonymous):

um.no?

OpenStudy (mathstudent55):

Ok, then you need to use the absolute value. Every time you see sqrt(c^2), you can replace it by |c|. When you see sqrt(c^4), you can replace it by c^2.

OpenStudy (mathstudent55):

\(\sqrt {16c^4} - \sqrt{ c^2}+3 \sqrt {c^2}+ \sqrt {9c^2}\) \(= 4c^2 - |c| + 3|c| + |3c| \) \(= 4c^2 - |c| + 3|c| + 3|c| \) Terms with |c| are like terms and can be combined together.

OpenStudy (anonymous):

ok..but I am still very confused on how to solve this problem

OpenStudy (anonymous):

sorry

OpenStudy (mathstudent55):

\(= 4c^2 +5 |c| \)

OpenStudy (anonymous):

ohh! I see how you kinda did the subsition method ;) thank you for helping me with this! I couldn't have done with out you! line no joke!

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (anonymous):

If I need help again.. (I don't right now) would you be willing to help me?

OpenStudy (anonymous):

im sorry but could you help me again on this: which factor would you cancel from the numerator and denominator to simplify x^2-3x-4 over x^2-16

OpenStudy (mathstudent55):

\( \dfrac{x^2 - 3x - 4}{x^2 - 16} \) You need to factor the numerator and the denominator. \( \dfrac{(x - 4)(x + 1)}{(x + 4)(x - 4) } \) Now you can divide the numerator and denominator by the same factor, x - 4.